OpenStudy (anonymous):

what is the solution to the rational equation 5/(x^2-3x+2) + 1/(x-2) = 1/3x-3

OpenStudy (compassionate):

Write it using the equation editor. You're very unclear with this.

OpenStudy (anonymous):

(5x^2-3x+2 ) plus (1/x-2) =(1/3x-2)

OpenStudy (anonymous):

5/x^2

OpenStudy (compassionate):

It cannot be solved.

OpenStudy (anonymous):

it can be i have 4 different answers to pick from x=5/2, x=-8 ,x=-7 , or x=9 I just dont know how to solve it.

OpenStudy (compassionate):

You're going to make me get my pen and and solve this. *Sigh* Alright give me a second.

OpenStudy (anonymous):

thank you

OpenStudy (compassionate):

One sec

OpenStudy (compassionate):

Making the final adjustments

OpenStudy (compassionate):

5(3x-3) + 1(x-1)(3x-3) = 1(x-1)(x-2) 15x - 15 + 3x^2 - 6x + 3 = x^2 - 3x + 2 x = -7 I even showed work. You're welcome. Can I have best answer?

OpenStudy (anonymous):

thank you so much!!!

OpenStudy (compassionate):

Welcome.